if an inelastic collision of two carts on a frictionless track collides, what is the momentum of cart A and B, before and after the collision? cart A has a mass of .5025 kg and cart B has a mass of .5491 kg. carts A initial velocity is .5595 m/s then collides and sticks to cart B, which was initially at rest, with a final velocity of .1622 m/s.
Answer #1

here,

the mass of cart A , mA = 0.5025 kg

mass of cart B , mB = 0.5491 kg

the initial velocity of A , uA = 0.5595 m/s

the initial momentum of cart A , Pai = mA * uA

Pai = 0.5025 * 0.5595 kg.m/s = 0.281 kg.m/s

the initial momentum of cart B, Pbi = mB * uB

Pbi = 0.5491 * 0 = 0 kg.m/s

the final velocity of B , vB = 0.1622 m/s

the final momentum of cart B, Pbf = mB * vB

Pbf = 0.5491 * 0.1622 = 0.0891 kg.m/s

let the final momentum of A be Paf

using conservation of momentum

Pai + Pbi = Paf + Pbf

0.281 + 0 = Paf + 0.0891

Paf = 0.19 kg.m/s